# GNN-Transformer Cooperative Architecture for Trustworthy Graph Contrastive Learning

### v1

@@ Introduction
Compared with traditional supervised learning, self-supervised learning eliminates the dependence on labels. As one of the most representative methods of self-supervised learning, contrastive learning has been widely applied in the graph domain, i.e., Graph Contrastive Learning (GCL). GCL utilizes augmentation strategies to generate multiple views and positive/negative pairs for node/graph representations (Zhu et al. 2020; Peng et al. 2020; Zhu, Sun, and Koniusz 2021; Yin et al. 2022; Liu et al. 2024). Proper augmentation strategies, graph encoders and loss functions are critical to the performance.
@@
Most of the existing GCL methods exploit GNNs as graph encoders (Velickovic et al. 2019; Hassani and Khasahmadi 2020; Mo et al. 2022; Lee, Lee, and Park 2022; Shen et al. 2023). GNNs iteratively propagate, transform, and aggregate representations from topological neighbors to progressively update node features (Kipf and Welling 2017; Hamilton, Ying, and Leskovec 2017; Zhang et al. 2022a). This process allows each node to capture structural information from its local neighborhood and gradually learn richer representations. With multi-layer stacking, these local features are progressively integrated to form a comprehensive perception of the entire graph structure. While this mechanism of GNNs is particularly effective in handling local dependencies, especially in capturing both direct and indirect relation between nodes, it fails to capture complex global structures and long-range dependencies within graphs (Xu et al. 2019; Garg, Jegelka, and Jaakkola 2020). In contrast, Graph Transformers (GTs) are capable of effectively capturing global dependencies between nodes with self-attention mechanisms (Veličković et al. 2017; Yun et al. 2019; Kreuzer et al. 2021). GTs can handle long-range dependencies within graphs, rather than being confined to local neighborhoods. By calculating attention scores between nodes, GTs update node representations with all other nodes in the graph, thereby effectively integrating global information. This mechanism enables GTs to model complex interactions and long-term dependencies between nodes. Therefore, they are more flexible and accurate in handling dependencies between distant nodes (Yun et al. 2019), which is ignored by existing GCL methods in choosing graph encoders to learn node/graph representations. However, the quadratic complexity of GTs limits the scalability. Currently, some works introduce linear attention mechanisms to improve scalability for large-scale graphs (Wu et al. 2022, 2024; Liang, Chen, and Liang 2024).
@@
Existing GCL methods mostly use InfoNCE as their loss functions (Zhu et al. 2020; Zhang et al. 2022b, 2023a), where each anchor point has only one positive pair, whereas all other across-view nodes are treated as negative pairs. Therefore, when calculating the loss, nodes from the same class as the anchor point are pushed away. This may lead to ignorance of the potential information from similar nodes when handling node feature similarity tasks (Shen et al. 2023). While this design effectively encourages the model to distinguish between positive and negative pairs, it may lead to insufficient aggregation of nodes from the same class, thereby affecting the ability to represent graph data and capture graph structures.
@@
In order to address the issues above, we propose a novel method called GNN-Transformer Cooperative Architecture for Trustworthy Graph Contrastive Learning (GTCA). GTCA utilizes GNN and NodeFormer, a linear GT as graph encoders to generate node representation views without random augmentation strategies. In addition, it utilizes topological property of graphs to generate topology structure view. Furthermore, we design a novel loss function to exploit the intersection of the node representation views and topology structure view as positive pairs. The contributions are summarized follows:
@@
  • We utilize both GNN and NodeFormer as graph encoders to capture comprehensive perception of graphs, which has not been well explored in the field of GCL to our best knowledge.
@@
Graph Contrastive Learning. GCL can be categorized into three mainstream paradigms: DGI framework (Velickovic et al. 2019), InfoNCE framework (Oord, Li, and Vinyals 2018) and BGRL framework (Thakoor et al. 2022).
Inspired by Deep InfoMax (Hjelm et al. 2019), DGI learns node representations by maximizing the mutual information between global and local features from the graph. MVGRL (Hassani and Khasahmadi 2020) extends DGI with the incorporation of graph diffusion with cross-view strategies, driving further advancement in unsupervised GCL. In spite of the competitive performance, global features in these methods may be insufficient to retain node-level embedding information.
Drawing on the significant success of InfoNCE, GRACE (Zhu et al. 2020) learns node representations by pulling together the same node across two augmented views and pushing apart different nodes. This principle is widely used in both node-level and graph-level representation learning to perform node classification (Zhu et al. 2021; Zhang et al. 2023b; Shen et al. 2023) and graph classification (You et al. 2020; Hassani and Khasahmadi 2020; Wei et al. 2023). While these methods achieve significant success on graphs, they still suffer from sampling bias issues.
To address the sampling bias problem, BGRL removes negative pairs with an exponential moving average method. AFGRL (Lee, Lee, and Park 2022) extends BGRL (Thakoor et al. 2022) with the elimination of graph augmentation and incorporating positive pairs sampling, which introduces sampling bias due to the limitation of random initialization in $k$-means.
Graph Augmentation. Recently, graph augmentation strategies play a crucial role in GCL. The existing graph augmentation strategies include node dropping (You et al. 2020), edge perturbation (Qiu et al. 2020; Zhang et al. 2023a), attribute masking (Zhu et al. 2021; Zhang et al. 2022b) and subgraph extraction (Hassani and Khasahmadi 2020). GRACE (Zhu et al. 2020) employs random edge perturbation and node feature masking to generate two views. GCA (Zhu et al. 2021) discards the randomness of augmentation and proposes an adaptive augmentation strategies which integrates both structural and attribute information. Similarly, GCS (Wei et al. 2023) utilizes the structural and semantic information of the graph to perform adaptive graph augmentation. Furthermore, CI-GCL (Tan et al. 2024) proposes a communityinvariant GCL framework to maintain graph community structure during learnable graph augmentation. Despite the considerable success, these methods still introduce biases due to they disturb the semantic information of the graph.
@@
Meanwhile, we utilize the NodeFormer (Wu et al. 2022) as the Transformer encoder, we can perform message passing using the following equation:
@@
Compared with existing graph-augmentation-based contrastive learning models, GTCA has several advantages. First, previous contrastive learning methods heavily rely on random or heuristic augmentation strategies, which may disturb the graph structure. On the contrary, GTCA does not require augmentation, making it adaptable to various graph datasets. Second, in contrast to existing GCL methods that utilize shared-weight GNNs, GTCA learns diverse graph feature representations with GNN and Transformer encoders to mitigate the issue of reduced feature diversity.
@@
According to (Oord, Li, and Vinyals 2018), the InfoNCE is a lower bound of the true Mutual Information (MI). Recent GCL methods use InfoNCE as the loss function (Zhu et al. 2020; Zhang et al. 2023a; Guo et al. 2023; Yu et al. 2024). Figure 3 gives an example of the InfoNCE loss. InfoNCE uses only one node to generate a positive pair of a given anchor node and utilizes all the other nodes to generate negative pairs. This may lead to ignorance of the potential information of similar nodes.
Figure 3: The red nodes denote the positive pair and the blue dotted lines with arrows are negative pairs formed with the anchor.
To address this issue, we design a novel contrastive loss function. Let $\bm{h}_{i}^{\theta}$ and $\bm{h}_{i}^{\varphi}$ denote the $\ell_{2}$-normalized GCN and NodeFormer embeddings of $v_{i}$ respectively. Select $\bm{h}_{i}^{\theta}$ as the anchor node, its positives come from three sources: (1) its NodeFormer embedding $\bm{h}_{i}^{\varphi}$; (2) GNN embeddings $\{\bm{h}_{j}^{\theta}\mid v_{j}\in\mathbf{P}_{i}\}$; (3) NodeFormer embeddings $\{\bm{h}_{j}^{\varphi}\mid v_{j}\in\mathbf{P}_{i}\}$. That is, the total number of positive pairs associated with node $v_{i}$ is $12|\mathbf{P}_{i}|+$ 1. Ultimately, the contrastive loss function is as follows:
@@
The minimization of Equation (7) will pull positive pairs closer and push negative pairs away. Since the two views are symmetric, we define the loss for the feature embedding $\bm{h}_{i}^{\varphi}$ corresponding to node $v_{i}$ in the other view similarly to Equation (7). The final loss function between the two views as follows:
@@
Given a graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, encoders $f_{\theta}$, $g_{\varphi}$ and a contrastive loss function $\mathcal{L}$ defined in Equation (11). $\mathbf{B}_{i}^{\theta}$, $\mathbf{B}_{i}^{\varphi}$, and $\mathbf{T}_{i}$ are $k$-NNs of node $v_{i}$, $i=1,2,\cdots,N$ under node representation views and topology structure view. Equation (5) generates the most trustworthy positive pairs for GCL.
@@
It is obvious that the intersection set contains the minimum number of positive pairs with specific $\mathbf{B}_{i}^{\theta}$, $\mathbf{B}_{i}^{\varphi}$, and $\mathbf{T}_{i}$. The numerator of Equation (7) sums the exponential calculation with respect to $\mathbf{P}_{i}$, whereas the denominator sums the exponential calculation with respect to both $\mathbf{P}_{i}$ and $\mathbf{N}_{i}$. Therefore, when calculating $\mathbf{P}_{i}$ with Equation (5), we get the lower bound of the numerator, and also upper bound of contrastive loss in Equation (7). When we minimize the upper bound of the loss in Equation (11), trustworthy GCL can be achieved. ∎
@@
The final output feature matrix $\bm{H}$ balances the contributions of $\bm{H}_{\theta}^{\prime}$ and $\bm{H}_{\varphi}^{\prime}$,which is applied to downstream node classification tasks. $\lambda$ is a tunable weight parameter. Algorithm 1 summarizes the learning process of GTCA.
@@
We compare GTCA with 12 state-of-the-art methods for semi-supervised node classification, including 2 semi-supervised GNNs, i.e., GCN (Kipf and Welling 2016) and GAT (Veličković et al. 2017), 2 semi-supervised GCL methods, i.e., CGPN (Wan et al. 2021b) and CG3 (Wan et al. 2021a), and 8 self-supervised GCL methods, i.e., DGI (Velickovic et al. 2019), GMI (Peng et al. 2020), MVGRL (Hassani and Khasahmadi 2020), GRACE (Zhu et al. 2020), GCA (Zhu et al. 2021), SUGRL (Mo et al. 2022), AFGRL (Lee, Lee, and Park 2022) and NCLA (Shen et al. 2023).
@@
We employ the self-supervised GCL baselines and GTCA to train and test an L2-regularized Logistic Regression (LR) classifier for node classification. For Cora dataset, we follows (Yang, Cohen, and Salakhudinov 2016) to randomly select 20 nodes per class for training, 500 nodes for validation, and the remaining nodes for testing. For Wiki-CS, Coauthor CS, Amazon-Computers and Amazon Photo datasets, we follows (Liu, Gao, and Ji 2020) to randomly select 20 nodes per class for training, 30 nodes per class for validation, and the remaining nodes for testing. It is noteworthy that for the self-supervised GCL baselines and GTCA model, we use the validation set to adjust the hyperparameters of the LR classifier. We perform 20 random splits of training, validation, and testing on each dataset and report the average performance of all algorithms across these splits. All experiments are implemented in PyTorch and conducted on a server with NVIDIA GeForce 3090 (24GB memory each). Table 2 shows the hyperparameter settings of GTCA on 5 datasets.
Datasets $k$ Embedding Dimension $E$ Weight Parameter $\lambda$ Learning Rate $lr$
@@
Table 2: Hyperparameter settings of GTCA on 5 datasets.
@@
Table 3: Node classification accuracy(%) comparison of different methods on 5 datasets.
Node Classification Results
Table 3 shows node classification accuracy on 5 benchmark graph datasets. On the whole, GTCA demonstrates superior performance across all 5 datasets. GTCA achieves state-of-the-art performance on Cora, Coauthor CS and Amazon-Photo datasets and ranks second on Wiki-CS and Amazon-Computers datasets.
The remarkable performance of GTCA can be attributed to two key aspects. First, we introduce an augmentation-free strategy for GCL, which avoids the potential risk of disturbing the underlying semantics of graphs. Second, GTCA employs a novel sampling strategy to generate multiple positive pairs within graphs, which contributes to the superior performance.
@@
Table 4 presents the ablation study results of GTCA and its variants without $\mathbf{T}_{i}$ and with different graph encoders including GNNs $f_{\theta},f_{\varphi}$ and NodeFormers $g_{\theta},g_{\varphi}$. We observe that there is an obvious decline in performance without topology structure view on 5 datasets. This indicates that topological information is essential for enhancing node classification accuracy. In terms of graph encoders, compared with GTCA, the use of either GNN-GNN or NodeFormer-NodeFormer results in inferior performance, which validates the reasonability of the module design.
@@
Hyperparameter Analysis
Figure 5 illustrates the node classification accuracy of GTCA on Cora and Amazon-Photo datasets with various hyperparameters, respectively. The optimal range of $k$ is between 300 and 600. It is obvious that increasing the embedding dimension $E$ generally leads to higher classification accuracy. It is also important to note that the choice of the hyperparameter $\lambda$ plays a crucial role in the performance. With the increase of $\lambda$, the accuracy of the node classification increases in the beginning, but starts to decrease slowly after reaching 0.7.
@@
Figure 5: Sensitivity analysis of hyperparameters $k$, $E$ and $\lambda$ on Cora and Amazon-Photo datasets.
@@
To provide a more intuitive presentation of the node embeddings, we utilize PCA (Abdi and Williams 2010) to visualize the node embeddings of DGI in Figure 4 (a), NCLA in Figure 4 (b) and GTCA in Figure 4 (c). Different colors represent different categories. Compared with DGI and NCLA, GTCA can effectively distinguish different classes of nodes.
Graph Encoders Analysis
Figure 6 shows that with a single graph encoder, correct ratio of positive pairs is lower than 40%. With both GCN and NodeFormer as graph encoders, the correct ratio of positive pairs varies from 30% to 60%. When GTCA exploits intersection of the $k$-NN neighborhood of GCN, Nodeformer node representation and topology structure views, the correct ratio of positive pairs exceed 80%. This highlights the necessity and effectiveness of the current graph encoder design.

### v2

@@ Introduction
Compared with traditional supervised learning, self-supervised learning eliminates the dependence on labels. As one of the most representative methods of self-supervised learning, contrastive learning has been widely applied in the graph domain, i.e., Graph Contrastive Learning (GCL). GCL utilizes multiple views and positive/negative pairs for node/graph representations (Zhu et al. 2020; Peng et al. 2020; Zhu, Sun, and Koniusz 2021; Yin et al. 2022; Liu et al. 2024), the performance of which are influenced by augmentation strategies, graph encoders and loss functions.
@@
Most of the existing GCL methods exploit GNNs as graph encoders (Velickovic et al. 2019; Hassani and Khasahmadi 2020; Mo et al. 2022; Lee, Lee, and Park 2022; Shen et al. 2023). GNNs iteratively propagate, transform, and aggregate representations from topological neighbors to progressively update node features (Kipf and Welling 2017; Hamilton, Ying, and Leskovec 2017; Zhang et al. 2022a). With multi-layer stacking, these local features are progressively integrated to form a comprehensive perception of the entire graph structure. While this mechanism is particularly effective in handling local dependencies, especially in capturing both direct and indirect relation between nodes, it fails to capture complex global structures and long-range dependencies within graphs (Xu et al. 2019; Garg, Jegelka, and Jaakkola 2020). In contrast, Graph Transformers (GTs) are capable of effectively capturing global dependencies between nodes with self-attention mechanisms (Veličković et al. 2017; Yun et al. 2019; Kreuzer et al. 2021). GTs can handle long-range dependencies within graphs, rather than being confined to local neighborhoods. By calculating attention scores between nodes, GTs update node representations with all other nodes in the graph, thereby effectively integrating global information. This mechanism enables GTs to model complex interactions and long-term dependencies between nodes. Therefore, they are more flexible and accurate in handling dependencies between distant nodes (Yun et al. 2019), which is ignored by existing GCL methods in choosing graph encoders to learn node/graph representations. However, the quadratic complexity of GTs limits the scalability. Currently, some works introduce linear attention mechanisms to improve scalability for large-scale graphs (Wu et al. 2022, 2024; Liang, Chen, and Liang 2024).
@@
Existing GCL methods mostly use InfoNCE as their loss functions (Zhu et al. 2020; Zhang et al. 2022b, 2023a), where each anchor point has only one positive pair. Therefore, when calculating the loss, nodes from the same class as the anchor point are pushed away. This may lead to ignorance of the potential information from similar nodes when handling node feature similarity tasks (Shen et al. 2023). While this design effectively encourages the model to distinguish between positive and negative pairs, it may lead to insufficient aggregation of nodes from the same class, thereby affecting the ability to represent graph data and capture graph structures.
@@
In order to address the issues above, we propose a novel method called GNN-Transformer Cooperative Architecture for Trustworthy Graph Contrastive Learning (GTCA). GTCA utilizes GCN and NodeFormer, a linear GT as graph encoders to generate node representation views without random augmentation strategies. In addition, it utilizes topological property of graphs to generate topology structure view. Furthermore, we design a novel loss function to exploit the intersection of the node representation views and topology structure view as positive pairs. The contributions are summarized follows:
@@
  • We utilize both GCN and NodeFormer as graph encoders to capture comprehensive perception of graphs, which has not been well explored in the field of GCL to our best knowledge.
@@
Graph Contrastive Learning. Current GCL methods can be categorized into three mainstream paradigms: DGI framework (Velickovic et al. 2019), InfoNCE framework (Oord, Li, and Vinyals 2018) and BGRL framework (Thakoor et al. 2022).
Inspired by Deep InfoMax (Hjelm et al. 2019), DGI learns node representations by maximizing the mutual information between global and local features from the graph. MVGRL (Hassani and Khasahmadi 2020) extends DGI with graph diffusion, driving further advancement in unsupervised GCL. In spite of the competitive performance, global features in these methods may be insufficient to retain node-level embedding information.
Inspired by the significant success of InfoNCE, GRACE (Zhu et al. 2020) learns node representations by pulling the same node together and pushing different nodes apart under two augmented views. This principle is widely used in both node classification (Zhu et al. 2021; Zhang et al. 2023b; Shen et al. 2023) and graph classification (You et al. 2020; Hassani and Khasahmadi 2020; Wei et al. 2023) tasks. While these methods achieve significant success, they still suffer from sampling bias issues.
To address the sampling bias problem, BGRL removes negative pairs with an exponential moving average method. AFGRL (Lee, Lee, and Park 2022) extends BGRL (Thakoor et al. 2022) with the augmentation-free strategy and node sampling technique. However, owning to the randomness of $k$-means, the sampling bias is introduced.
Graph Augmentation. The existing graph augmentation strategies include node dropping (You et al. 2020), edge perturbation (Qiu et al. 2020; Zhang et al. 2023a), attribute masking (Zhu et al. 2021; Zhang et al. 2022b) and subgraph extraction (Hassani and Khasahmadi 2020). GRACE (Zhu et al. 2020) employs random edge perturbation and node feature masking to generate two views. GCA (Zhu et al. 2021) proposes an adaptive augmentation strategy which integrates both structural and attribute information. Similarly, GCS (Wei et al. 2023) utilizes the structural and semantic information to achieve adaptive graph augmentation. Furthermore, CI-GCL (Tan et al. 2024) proposes a community invariant GCL framework to maintain graph community structure. Despite the considerable success, bias is introduced due to the disturbance of the semantic.
@@
Meanwhile, we utilize the NodeFormer (Wu et al. 2022) as the Transformer encoder with the following equation:
@@
Compared with existing augmentation based contrastive learning models, GTCA has several advantages. First, GTCA discards random or heuristic augmentation strategies which may disturb the graph structure. Second, in contrast to existing GCL methods that utilize shared-weight GNNs, GTCA uses GNN and Transformer encoders to learn diverse graph feature representations.
@@
According to (Oord, Li, and Vinyals 2018), the InfoNCE is a lower bound of the true Mutual Information (MI). Recent GCL methods use InfoNCE as the loss function (Zhu et al. 2020; Zhang et al. 2023a; Guo et al. 2023; Yu et al. 2024). Figure 3 gives an example of the InfoNCE loss. For a given anchor node, InfoNCE uses only one node to generate a positive pair and utilizes all the other nodes to generate negative pairs. This may lead to ignorance of the potential information of similar nodes.
Figure 3: The red nodes denote the positive pair and the blue dotted lines with arrows are negative pairs of the anchor.
To address this issue, we design a novel contrastive loss function. Let $\bm{h}_{i}^{\theta}$ and $\bm{h}_{i}^{\varphi}$ denote the $\ell_{2}$-normalized GCN and NodeFormer embeddings of $v_{i}$ respectively. Select $\bm{h}_{i}^{\theta}$ as the anchor node, its positives come from three sources: (1) its NodeFormer embedding $\bm{h}_{i}^{\varphi}$; (2) GCN embeddings $\{\bm{h}_{j}^{\theta}\mid v_{j}\in\mathbf{P}_{i}\}$; (3) NodeFormer embeddings $\{\bm{h}_{j}^{\varphi}\mid v_{j}\in\mathbf{P}_{i}\}$. That is, the total number of positive pairs associated with node $v_{i}$ is $12|\mathbf{P}_{i}|+$ 1. Ultimately, the contrastive loss function is as follows:
@@
The minimization of Equation (7) will pull positive pairs closer and push negative pairs away. Since the two views are symmetric, we define the loss for the feature embedding $\bm{h}_{i}^{\varphi}$ corresponding to node $v_{i}$ in the other view similarly to Equation (7). Then, the final loss function is as follows:
@@
Given a graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, encoders $f_{\theta}$, $g_{\varphi}$ and a contrastive loss function $\mathcal{L}$ defined in Equation (11). $\mathbf{B}_{i}^{\theta}$, $\mathbf{B}_{i}^{\varphi}$, and $\mathbf{T}_{i}$ are $k$-NNs of node $v_{i}$, $i=1,2,\cdots,N$ under two node representation views and one topology structure view. Equation (5) generates the most trustworthy positive pairs for GCL.
@@
It is obvious that the intersection set contains the minimum number of positive pairs with specific $\mathbf{B}_{i}^{\theta}$, $\mathbf{B}_{i}^{\varphi}$, and $\mathbf{T}_{i}$. The numerator of Equation (7) sums the exponential calculation with respect to $\mathbf{P}_{i}$, whereas the denominator sums the exponential calculation with respect to both $\mathbf{P}_{i}$ and $\mathbf{N}_{i}$. Therefore, when calculating $\mathbf{P}_{i}$ with Equation (5), we get the lower bound of the numerator, and also the upper bound of the contrastive loss in Equation (7). When we minimize the upper bound of the loss in Equation (11), trustworthy GCL can be achieved. ∎
@@
The final output feature matrix $\bm{H}$, a linear combination of $\bm{H}_{\theta}^{\prime}$ and $\bm{H}_{\varphi}^{\prime}$, is applied to downstream node classification tasks. $\lambda$ is a tunable weight parameter. Algorithm 1 summarizes the overall procedure of GTCA.
@@
We compare GTCA with 12 state-of-the-art methods, including 2 semi-supervised GNNs, i.e., GCN (Kipf and Welling 2016) and GAT (Veličković et al. 2017), 2 semi-supervised GCL methods, i.e., CGPN (Wan et al. 2021b) and CG3 (Wan et al. 2021a), and 8 self-supervised GCL methods, i.e., DGI (Velickovic et al. 2019), GMI (Peng et al. 2020), MVGRL (Hassani and Khasahmadi 2020), GRACE (Zhu et al. 2020), GCA (Zhu et al. 2021), SUGRL (Mo et al. 2022), AFGRL (Lee, Lee, and Park 2022) and NCLA (Shen et al. 2023).
@@
We employ the above methods for node classification. For Cora dataset, we follows (Yang, Cohen, and Salakhudinov 2016) to randomly select 20 nodes per class for training, 500 nodes for validation, and the remaining nodes for testing. For Wiki-CS, Coauthor CS, Amazon-Computers and Amazon Photo datasets, we follows (Liu, Gao, and Ji 2020) to randomly select 20 nodes per class for training, 30 nodes per class for validation, and the remaining nodes for testing. We perform 20 random splits of training, validation, and testing on each dataset and report the average performance of all algorithms across these splits. All experiments are implemented in PyTorch and conducted on a server with NVIDIA GeForce 3090 (24GB memory each). Table 2 shows the hyperparameter settings of GTCA on 5 datasets.
Datasets $k$ $E$ $\lambda$ $lr$
@@
Table 2: Hyperparameter settings of GTCA on 5 datasets. $lr$ is the learning rate.
Node Classification Results
Table 3 shows node classification accuracy on 5 benchmark graph datasets. On the whole, GTCA demonstrates superior performance across all 5 datasets. GTCA ranks first on three datasets and ranks second on the other two datasets.
The remarkable performance of GTCA can be attributed to two key aspects. First, GTCA introduces an augmentation-free strategy, which avoids the potential risk of disturbing the underlying semantics of graphs. Second, GTCA employs a novel sampling strategy to generate trustworthy positive pairs and negative pairs.
@@
Table 3: Node classification accuracy (%) comparison on 5 datasets.
@@
Table 4 presents the ablation study results of GTCA and its variants without $\mathbf{T}_{i}$ and with different graph encoders including GNNs $f_{\theta},f_{\varphi}$ and NodeFormers $g_{\theta},g_{\varphi}$. We observe that there is an obvious decline in performance without topology structure view on 5 datasets. This indicates that topological information is essential for enhancing node classification accuracy. In terms of graph encoders, compared with GTCA, the use of either GCN-GCN or NodeFormer-NodeFormer results in inferior performance, which validates the reasonability of the module design.
@@
Graph Encoders Analysis
Figure 4 shows that with a single graph encoder, correct ratio of positive pairs is lower than 40%. With both GCN and NodeFormer as graph encoders, the correct ratio of positive pairs varies from 30% to 60%. When GTCA exploits intersection of the $k$-NN neighborhood of GCN, Nodeformer node representation and topology structure views, the correct ratio of positive pairs exceeds 80%. This highlights the necessity and effectiveness of the current graph encoder design.
Sensitivity Analysis
Figure 5 illustrates the node classification accuracy of GTCA on Cora and Amazon-Photo datasets with various hyperparameters, respectively. The optimal range of $k$ is between 300 and 600. When $k$ is too small, GTCA has inferior performance due to a scarcity of positive pairs. Conversely, a large $k$ introduces too much redundant information. Thus, it is necessary to search a proper $k$. In addition, increasing embedding dimension $E$ improve classification accuracy. Moreover, the hyperparameter $\lambda$ is crucial for model performance. Initially, an increase in $\lambda$ improves node classification accuracy. However, when $\lambda$ exceeds 0.7, the accuracy will decrease.
Figure 5: Accuracy vs. hyperparameters $k$, $E$ and $\lambda$ on Cora and Amazon-Photo datasets.
@@
To provide a more intuitive presentation of the node embeddings, we utilize PCA (Abdi and Williams 2010) to visualize the node embeddings of DGI in Figure 6 (a), NCLA in Figure 6 (b) and GTCA in Figure 6 (c). Different colors represent different categories. Compared with DGI and NCLA, GTCA can distinguish different classes of nodes much more effectively.